7(a+b)=b-a
4(3a+2b)=b-8

Answer by Alan3354(69443)

7(a+b)=b-a
4(3a+2b)=b-8
-----------------
If you want to solve for a & b, you should say that.
---
7(a+b)=b-a --> 7a + 7b = b - a ---> 8a = -6b ---> a = -3b/4
=====================
4(3a+2b)=b-8 --> 12a + 8b = b-8
Sub for a
12*(-3b/4) + 8b = b-8
You can do the rest